\begin{thebibliography}{10}

\bibitem{MS8977}
Gregory Moore and Nathan Seiberg.
\newblock Classical and quantum conformal field theory.
\newblock {\em Commun. Math. Phys.}, 123(2):177--254, June 1989.

\bibitem{W8951}
Edward Witten.
\newblock Quantum field theory and the {Jones} polynomial.
\newblock {\em Commun. Math. Phys.}, 121(3):351--399, September 1989.

\bibitem{W8987}
Xiao-Gang Wen.
\newblock Vacuum degeneracy of chiral spin states in compactified space.
\newblock {\em Phys. Rev. B}, 40(10):7387--7390, October 1989.

\bibitem{W150605768}
Xiao-Gang Wen.
\newblock A theory of {2+1D} bosonic topological orders.
\newblock {\em Nat. Sci. Rev.}, 3(1):68--106, November 2015.

\bibitem{KZ150201690}
Liang Kong, Xiao-Gang Wen, and Hao Zheng.
\newblock Boundary-bulk relation for topological orders as the functor mapping
  higher categories to their centers.
\newblock {\em arXiv:1502.01690}, 2015.

\bibitem{TW191202817}
Ryan {Thorngren} and Yifan {Wang}.
\newblock {Fusion Category Symmetry I: Anomaly In-Flow and Gapped Phases}.
\newblock {\em J. High Energ. Phys.}, 2024:132, December 2024.

\bibitem{KZ200514178}
Liang {Kong}, Tian {Lan}, Xiao-Gang {Wen}, Zhi-Hao {Zhang}, and Hao {Zheng}.
\newblock {Algebraic higher symmetry and categorical symmetry: A holographic
  and entanglement view of symmetry}.
\newblock {\em Phys. Rev. Res.}, 2(4):043086, October 2020.

\bibitem{Wen1}
Xiao-Gang Wen.
\newblock A theory of 2+1d bosonic topological orders.
\newblock {\em Nat. Sci. Rev.}, 3(1):68--106, 2015.

\bibitem{BNRW}
Paul Bruillard, Siu-Hung Ng, Eric~C. Rowell, and Zhenghan Wang.
\newblock On classification of modular categories by rank.
\newblock {\em Int. Math. Res. Not. IMRN}, (24):7546--7588, 2016.

\bibitem{NRWW}
Siu-Hung Ng, Eric~C. Rowell, Zhenghan Wang, and Xiao-Gang Wen.
\newblock Reconstruction of modular data from {${\rm SL}_2(\Bbb Z)$}
  representations.
\newblock {\em Comm. Math. Phys.}, 402(3):2465--2545, 2023.

\bibitem{NRW}
Siu-Hung Ng, Eric~C. Rowell, and Xiao-Gang Wen.
\newblock Classification of modular data up to rank 11.
\newblock {\em arXiv:2308.09670}, 2023.

\bibitem{MS1}
Micha\"el Mignard and Peter Schauenburg.
\newblock Modular categories are not determined by their modular data.
\newblock {\em Lett. Math. Phys.}, 111(3):Paper No. 60, 9, 2021.

\bibitem{BPD}
Parsa Bonderson, Colleen Delaney, C\'{e}sar Galindo, Eric~C. Rowell, Alan Tran,
  and Zhenghan Wang.
\newblock On invariants of modular categories beyond modular data.
\newblock {\em J. Pure Appl. Algebra}, 223(9):4065--4088, 2019.

\bibitem{KMS}
Ajinkya Kulkarni, Micha\"el Mignard, and Peter Schauenburg.
\newblock A topological invariant for modular fusion categories.
\newblock {\em arXiv:1806.03158}, 2023.

\bibitem{EGNO}
Pavel Etingof, Shlomo Gelaki, Dmitri Nikshych, and Victor Ostrik.
\newblock {\em Tensor categories}, volume 205 of {\em Mathematical Surveys and
  Monographs}.
\newblock American Mathematical Society, Providence, RI, 2015.

\bibitem{BaKi}
Bojko Bakalov and Alexander Kirillov, Jr.
\newblock {\em Lectures on tensor categories and modular functors}, volume~21
  of {\em University Lecture Series}.
\newblock American Mathematical Society, Providence, RI, 2001.

\bibitem{KasBook}
Christian Kassel.
\newblock {\em Quantum groups}, volume 155 of {\em Graduate Texts in
  Mathematics}.
\newblock Springer-Verlag, New York, 1995.

\bibitem{NS1}
Siu-Hung Ng and Peter Schauenburg.
\newblock {Higher Frobenius-Schur Indicators for Pivotal Categories}.
\newblock In {\em {Hopf Algebras and Generalizations}}, volume 441 of {\em
  Contemp. Math.}, pages 63--90. Amer. Math. Soc., Providence, RI, 2007.

\bibitem{NS2}
Siu-Hung Ng and Peter Schauenburg.
\newblock Central invariants and higher indicators for semisimple quasi-{H}opf
  algebras.
\newblock {\em Trans. Amer. Math. Soc.}, 360(4):1839--1860, 2008.

\bibitem{NS4}
Siu-Hung Ng and Peter Schauenburg.
\newblock Congruence subgroups and generalized {F}robenius-{S}chur indicators.
\newblock {\em Comm. Math. Phys.}, 300(1):1--46, 2010.

\bibitem{PSS}
G.~Pradisi, A.~Sagnotti, and Ya.~S. Stanev.
\newblock Planar duality in {${\rm SU}(2)$} {WZW} models.
\newblock {\em Phys. Lett. B}, 354(3-4):279--286, 1995.

\bibitem{BHS}
L.~Borisov, M.~B. Halpern, and C.~Schweigert.
\newblock Systematic approach to cyclic orbifolds.
\newblock {\em Internat. J. Modern Phys. A}, 13(1):125--168, 1998.

\bibitem{BCT}
Daniel Barter, Corey Jones, and Henry Tucker.
\newblock Eigenvalues of rotations and braids in spherical fusion categories.
\newblock {\em J. Algebra}, 515:52--76, 2018.

\bibitem{Bantay97}
Peter Bantay.
\newblock The {F}robenius-{S}chur indicator in conformal field theory.
\newblock {\em Phys. Lett. B}, 394(1-2):87--88, 1997.

\end{thebibliography}

\documentclass{amsart}
\usepackage[all, knot, color]{xy}
\usepackage[utf8]{inputenc}
\usepackage{amsmath, latexsym, amssymb, fancyhdr, color, mathtools}
\usepackage{marginnote}
\usepackage{bbold}
\usepackage{changepage}
\usepackage{amscd}
\usepackage{bm}
\usepackage[bbgreekl]{mathbbol}
\usepackage{tikz}
\usetikzlibrary{cd, calc, knots}

\usepackage{graphicx}
\def\bx{{\bm{x}}}
\def\bX{{\bm{X}}}
\def\a{{\alpha}}
\def\b{{\beta}}

\def\e{{\epsilon}}
\def\w{{\omega}}

\def\o{{\otimes}}
\def\s{{\sigma}}

\def\vec{{\operatorname{Vec}}}

\def\SL2{{\mbox{SL}_2(\BZ)}}
\def\Mp2{{\mbox{Mp}_2(\BZ)}}

\def\1{{\mathbb{1}}}
\def\to{\rightarrow}

\def\BC{{\mathbb{C}}}

\def\BZ{{\mathbb{Z}}}

\def\CC{\mathcal{C}}

\def\DD{\mathcal{D}}

\def\lcm{{\operatorname{lcm}}}

\newcommand\Irr{{\operatorname{Irr}}}

\def\id{{\operatorname{id}}}
\def\lcm{{\operatorname{lcm}}}

\def\Tr{{\operatorname{Tr}}}

\def\ord{{\operatorname{ord}\,}}

\def\ev{{\operatorname{ev}}}
\def\coev{{\operatorname{coev}}}
\newcommand{\red}[1]{{\color{red} #1}}

\newcommand\ol[1]{\overline{#1}}

\newtheorem{thm}{Theorem}[section]

\newtheorem{prop}[thm]{Proposition}

\theoremstyle{definition}

\newtheorem{remark}[thm]{Remark}

\usetikzlibrary{decorations.markings}
\tikzset{->-/.style={decoration={
			markings,
			mark=at position #1 with {\arrow{>}}},postaction={decorate}}}
\tikzset{-<-/.style={decoration={
			markings,
			mark=at position #1 with {\arrow{<}}},postaction={decorate}}}

\title{Recovering $R$-symbols from modular data}

\begin{document}
\author{Siu-Hung Ng}
\address{Department of Mathematics,
Louisiana State University,
Baton Rouge, LA 70803, USA}
\email[Siu-Hung~Ng]{rng@@math.lsu.edu}
\author{Eric C. Rowell}
\address{ Department of Mathematics, Texas A\&M University, College
Station, TX 77843, USA }
\email[Eric~C.~Rowell]{rowell@math.tamu.edu}
\author{Xiao-Gang Wen}
\address{Department of Physics, Massachusetts Institute of Technology,
Cambridge, Massachusetts 02139, USA}
\email[Xiao-Gang~Wen]{xgwen@mit.edu}

\begin{abstract}
 Given a premodular category $\CC$, we show that its $R$-symbol can be recovered from its $T$-matrice,  fusion coefficients and some 2nd generalized Frobenius-Schur indicators. In particular, if $\CC$ is modular, its $R$-symbols for a certain gauge choice are completely determined by its modular data. 
\end{abstract}
\maketitle

\section{Introduction} 

Modular tensor categories (MTCs) are spherical nondegenerate braided fusion
categories. These categories naturally arise in both mathematics and physics
such that  the theory of subfactors, rational conformal theory \cite{MS8977},
topological quantum field theory \cite{W8951}, topological orders
\cite{W8987,W150605768}, and many more. Recently, MTC theory becomes more
important due to a realization that symmetries in 1-dimensional quantum systems
are actually described by MTCs in trial Witt class
\cite{KZ150201690,TW191202817,KZ200514178}, rather than by groups that can
only describe a subset of symmetries.

Associated to an MTC $\CC$ is a pair $(S,T)$ of complex square matrices, called the \emph{modular data} of $\CC$, which are  indexed by a complete set $\Irr(\CC)$ of simple objects of $\CC$, and uniquely determined by $\CC$ up to some permutations on $\Irr(\CC)$. The number $|\Irr(\CC)|$ is called the \emph{rank} of the modular data $(S,T)$.

There has been numerous efforts to generate modular data of small ranks (cf. \cite{Wen1, BNRW, NRWW, NRW}). However, it has been shown by examples that modular data of an MTC does not determine the underlying MTC  \cite{MS1}. More precisely, the nonabelian group $G$ of order $55$ admits two non-cohomologous 3-cocycles $\w$ and $\w'$ such that the centers $Z(\vec_G^\w)$ and  $Z(\vec_G^{\w'})$ are inequivalent braided fusion categories but their modular data are the same, up to a permutation. 

To completely determine an MTC $\CC$, up to equivalence, one may need its associativity constraints and braidings. It is natural to ask whether there is any invariant of MTCs beyond modular data without knowing the complete information of its associativity constraint and braiding. The Whitehead links \cite{BPD} and the Borromean links \cite{KMS} colored by simple objects of an MTC were shown to be invariants beyond modular data, and they can distinguish $Z(\vec_G^{\w})$ for any 3-cocycle $\w$ of the nonabelian group $G$ of order 55.

These invariants motivate us to consider the braiding or the gauge equivalent classes of $R$-symbols of an MTC as invariants beyond modular data. However, we find that, for some \emph{gauge choice}, the $R$-symbols can be determined by its modular data of any MTC. 

In general, $R$-symbols of a premodular category $\CC$ can determined by its $T$-matrix, fusion rules and the second generalized Frobenius-Schur indicators.  When $\CC$ is modular, its fusion rules and generalized Frobenius-Schur indicators can be expressed in terms of its modular data. 

The article assumes some basic knowledge on modular tensor categories, graphical calculus and Frobenius-Schur indicators for pivotal categories. The readers are referred to \cite{EGNO, BaKi, KasBook} for more details on tensor categories and their graphical calculus. The categorical development of Frobenius-Schur indicators can be found in \cite{NS1, NS2, NS4}.
\section{$R$-symbols of premodular categories} \label{s1}
Let $\CC$ be a premodular (or ribbon) category  over $\BC$, and $\Irr(\CC)$ denote a complete set of nonisomorphic simple objects containing the unit object $\1$. 

For any objects $x,y \in \CC$, the morphism space of $\CC$ from $x$ to $y$, denoted by $\CC(x,y)$, is a finite-dimensional $\BC$-linear space.  Given any $a,b,c \in \Irr(\CC)$, one can choose a basis $
B^{a,b}_c$ for the $\BC$-linear space $\CC(a \o b, c)$. The collection 
$$
B= \big\{B^{a,b}_c \mid a,b,c \in \Irr(\CC)\big\}
$$
is simply called a \emph{gauge choice} of $\CC$. In particular, we write  $N^{a,b}_c$ for $|B^{a,b}_c|$. The semisimplicity of $\CC$ immediately implies  that $N_c^{a,b}$ is the fusion coefficient determined by the tensor product
$$
a \o b = \bigoplus_{c \in \Irr(\CC)} N_c^{a,b} \, c\quad \text{for any } a, b \in \Irr(\CC)\,.
$$

For the purpose of using graphical calculus, we may assume $\CC$ is a strict pivotal category without loss of generality \cite[tthm. 2.2]{NS1}. A morphism $\pi \in \CC(a \o b, c)$ will be depicted as a trivalent tree: 
$\pi=  
\xy 0;/r1pc/:
(0,0)*{\bullet},
(0,0); (-1,1.2)*+{a} **\dir{-},
(0,0); (1,1.2)*+{b} **\dir{-},
(0,0); (0,-1.2)*+{c} **\dir{-}
\endxy
$. 
We may label the vertex of the trivalent tree to distinguish  possibly different morphisms in $\CC(a \o b, c)$. 
Thus, we may write 
$$
B^{a,b}_c = \left\{ \xy 0;/r1pc/:
(0,0)*{\bullet},
(0,0); (-1,1.2)*+{a} **\dir{-},
(0,0); (1,1.2)*+{b} **\dir{-},
(0,0); (0,-1.2)*+{c} **\dir{-},
(0,.5)*{\scriptstyle \a}
\endxy
: \a=1,\dots,N^{a,b}_c 
\right\}\,.
$$

The braiding $R_{x,y}: x \o y \to y \o x$ of $\CC$ for any $x,y \in \CC$ is depicted by the diagram: 
$$
R_{x,y} \, = \, \vcenter{\xy 0;/r1.5pc/:
(0,1)\vtwistneg[1],
(0,1.2)*{\scriptstyle{x}},
(1,1.2)*{\scriptstyle{y}},
(0,-.3)*{\scriptstyle{y}},
(1,-.3)*{\scriptstyle{x}},
\endxy}\,.
$$
For any $x, y \in \CC$, the braiding $R_{x,y}$ acts as a linear isomorphism, 
$$
R_{x,y}^z: \CC(y \o x, z) \to \CC(x \o y, z)
$$
by composition, for any $z \in \CC$. Specifically, $R^z_{x,y}(\pi) = \pi \circ R_{x,y}$ for $\pi \in \CC(y \o x, z)$. In terms of diagrams, 
$$
R^z_{x,y}: \vcenter{\xy 0;/r1pc/:
(0,0)*{\bullet},
(0,0); (-1,1.2)*+{\scriptstyle{y}} **\dir{-},
(0,0); (1,1.2)*+{\scriptstyle{x}} **\dir{-},
(0,0); (0,-1.2)*+{\scriptstyle{z}} **\dir{-}
\endxy} \mapsto \,
\vcenter{\xy 0;/r1pc/:
\vtwistneg~{(-1,2)*+{\scriptstyle{x}}}{(0.5,2)*+{\scriptstyle{y}}}{(-1,.5)}{(0.5,.5)},
(-.25,-.3)*{\bullet},
(-.25,-.3); (-1,.5) **\crv{(-1,.2)},
(-.25,-.3); (.5,.5) **\crv{(.5, .2)},
(-.25,-.3); (-.25,-1.2)*+{\scriptstyle{z}} **\dir{-},
\endxy}\,.
$$
Since $R_{x,y}$ is an isomorphism in $\CC$, if $B^{y,x}_z$ is a basis for $\CC(y \o x, z)$ then 
$$
B^{x,y}_z=\{R_{x,y}(\pi) \mid \pi \in B^{y,x}_z\}
$$
is a basis for $\CC(x \o y, z)$. 

Let $\theta$ be the ribbon structure of $\CC$, which means  $\theta$ is a natural isomorphism of the identity functor of $\CC$ satisfying the conditions:
\begin{equation}\label{eq:ribbon}
    (\theta_x)^*  = \theta_{x^*} \quad\text{and}\quad \theta_{x \o y} = (\theta_x \o \theta_y) R_{y,x} R_{x,y}
\end{equation}
for any $x, y \in \CC$. Thus, for any $a \in \Irr(\CC)$, $\theta_a$ is a scalar multiple of $\id_a$, and we will abuse notation by denoting this scalar by $\theta_a$. In general, $\theta_x: x \to x$ is an isomorphism in $\CC$. Since $\CC$ is assumed to be strict pivotal, we can depict $\theta_x$ as
$$
\theta_x \,=\, \vcenter{\xy 0;/r1pc/:
(0,2)\vcap[1],
\vtwistneg~{(1,2)}{(2,2)}{(1, 1)}{(2,1)},
(2,2); (2,3) **\dir{-}?(1)+(0.4,0)*{\scriptstyle x},,
(2,1); (2,0) **\dir{-}?(1)+(0.4,0)*{\scriptstyle x},
(0,2); (0,1) **\dir{-},
(0,1)\vcap[-1],\endxy}\,,
$$
where the coevaluation $\coev: \1 \to y \o y^*$ and the evaluation $\ev: y^* \o y \to \1$ morphisms are respectively depicted by the diagrams  
$$
\vcenter{\xy 0;/r1pc/:
(0,2.1)\vcap[2],
(2.2,1.8)*{\scriptstyle y^*},
(0,1.8)*{\scriptstyle y},
(5,2)*{,},
(8,2.7)\vcap[-2],
(8,3.3)*{\scriptstyle y^*},
(10,3.3)*{\scriptstyle y},
(12,2)*{.},
\endxy}
$$
In particular, for $a, b \in \Irr(\CC)$,
$$
\vcenter{\xy 0;/r1pc/:
(0,2)\vcap[1],
\vtwistneg~{(1,2)}{(2,2)}{(1, 1)}{(2,1)},
(2,2); (2,3) **\dir{-}?(1)+(0.4,0)*{\scriptstyle a},,
(2,1); (2,0) **\dir{-}?(1)+(0.4,0)*{\scriptstyle a},
(0,2); (0,1) **\dir{-},
(0,1)\vcap[-1],\endxy} = \theta_a\, \,
\vcenter{\xy 0;/r1pc/:
(2,3); (2,0) **\dir{-}?(.5)+(0.4,0)*{\scriptstyle a},\endxy} \quad\text{ and so }
(\theta_a)^* = 
\vcenter{\xy 0;/r1pc/:
(0,2)\vcap[1],
\vtwistneg~{(1,2)}{(2,2)}{(1, 1)}{(2,1)},
(2,2); (2,3) **\dir{-}?(.8)+(-0.4,0)*{\scriptstyle a},
(2,3)\vcap[1],
(2,1); (2,0) **\dir{-},
(0,2); (0,1) **\dir{-},
(-.5,0)\vcap[-2.5],
(0,1)\vcap[-1],
(-0.5,0); (-.5,4) **\dir{-}?(1)+(-0.4,0)*{\scriptstyle a^*},
(3,3); (3,-2) **\dir{-}?(1)+(0.4,0)*{\scriptstyle a^*},
\endxy} = \theta_a\, \,
\vcenter{\xy 0;/r1pc/:
(2,3); (2,0) **\dir{-}?(.5)+(0.4,0)*{\scriptstyle a^*},\endxy}
$$
and the second part of \eqref{eq:ribbon} becomes
\begin{equation} \label{eq:ribbon2}
 \vcenter{\xy 0;/r1.5pc/:
\vtwistneg~{(-1,3.5)*+{\scriptstyle{a}}}{(0.5,3.5)*+{\scriptstyle{b}}}{(-1,2)}{(0.5,2)},
\vtwistneg~{(-1,2)}{(0.5,2)}{(-1,.5)*+{\scriptstyle{a}}}{(0.5,.5)*+{\scriptstyle{b}}},
\endxy} =
\theta_a^{-1}\theta_b^{-1}\, \vcenter{\xy 0;/r1pc/:
\vcross~{(0,1)}{(.5,1)}{(0,0)}{(.5, 0)},
\vcross~{(.5,0)}{(1.5,0)}{(.5,-1)}{(1.5,-1.5)},
\vcross~{(0,-1)}{(.5,-1)}{(0,-2)}{(.5,-2)},
\vcross~{(-.5,0)}{(0,0)}{(-.5,-1)}{(0,-1)},
(-.5,-1);(-.5,-2) **\dir{-},
(-.5,0);(-.5,1) **\dir{-},
(-1.5,1)\vcap[1],
(-2,1)\vcap[2],
(.5,1);(.5,2)*+{\scriptstyle a} **\dir{-},
(1.5,0);(1.5,2)*+{\scriptstyle b} **\dir{-},
(-1.5,-2)\vcap[-1],
(.5,-2);(.5,-3)*+{\scriptstyle a} **\dir{-},
(-1.5,1);(-1.5,-2)**\dir{-},
(-2,-2)\vcap[-2],
(-2,1);(-2,-2) **\dir{-},
(1.5,-1.5);(1.5,-3)*+{\scriptstyle{b}} **\dir{-},
\endxy}\,.
\end{equation}

\begin{prop} \label{p1}
Let $\CC$ be a premodular category over $\BC$ with the ribbon structure $\theta$ and braiding $R$. Assume an arbitrary complete order on $\Irr(\CC)$.  There exists a gauge choice $B$ of $\CC$ such that the matrix $[R_{a,b}^c]$ relative to the corresponding bases  of $\CC(b \o a, c)$ and $\CC(a \o b, c)$ is a diagonal matrix for any $a,b,c \in \Irr(\CC)$. Moreover,
$$
[R_{a,b}^c] = \left\{\begin{array}{ll}
\id & \text{if } a > b,\\
\frac{\theta_c}{\theta_a \theta_b} \id & \text{if } a < b,\\
\frac{\sqrt{\theta_c}}{\theta_a} \Delta_a^c & \text{if } a = b,
\end{array}\right.
$$
where $\theta_c$ is the scalar of the ribbon structure specified at $c \in \Irr(\CC)$, $\sqrt{\theta_c}$ is any chosen square root of $\theta_c$ for each $c \in \Irr(\CC)$, and $\Delta_a^c$ is a signed diagonal matrix.
\end{prop}
\begin{proof}
For $a, b, c \in \Irr(\CC)$ with $a < b$, we (arbitrarily) fix a basis $B^{a,b}_c$ for $\CC(a\o b, c)$. If $a > b$, then $B_c^{a,b}:=R_{a,b}^c( B^{b,a}_c)$ is a basis for $\CC(a\o b, c)$ since $R^c_{a,b}$ is a $\BC$-linear isomorphism. Thus, if $a > b$, the matrix $[R^c_{a,b}]$ of  $R^c_{a,b}$ relative to these bases is the identity matrix $\id$. On the other hand, if $a < b$, then 
$$
B_c^{b, a}=\left\{
\vcenter{\xy 0;/r1pc/:
\vtwistneg~{(-1,2)*+{\scriptstyle{b}}}{(0.5,2)*+{\scriptstyle{a}}}{(-1,.5)}{(0.5,.5)},
(-.25,-.3)*{\bullet},
(-.25,-.3); (-1,.5) **\crv{(-1,.2)},
(-.25,-.3); (.5,.5) **\crv{(.5, .2)},
(-.25,-.3); (-.25,-1.2)*+{\scriptstyle{c}} **\dir{-},
(-.25,.1)*{\scriptstyle{\a}},
\endxy}: 
\vcenter{\xy 0;/r1pc/:
(0,0)*{\bullet},
(0,0); (-1,1.2)*+{\scriptstyle{a}} **\dir{-},
(0,0); (1,1.2)*+{\scriptstyle{b}} **\dir{-},
(0,0); (0,-1.2)*+{\scriptstyle{c}} **\dir{-},
(0,.4)*{\scriptstyle{\a}},
\endxy} \in B_c^{a,b}\right\}
$$
and 
$$
R_{a,b}^c\left(
\vcenter{\xy 0;/r1pc/:
\vtwistneg~{(-1,2.5)*+{\scriptstyle{b}}}{(0.5,2.5)*+{\scriptstyle{a}}}{(-1,.5)}{(0.5,.5)},
(-.25,-.3)*{\bullet},
(-.25,-.3); (-1,.5) **\crv{(-1,.2)},
(-.25,-.3); (.5,.5) **\crv{(.5, .2)},
(-.25,-.3); (-.25,-1.2)*+{\scriptstyle{c}} **\dir{-},
(-.25,.1)*{\scriptstyle{\a}},
\endxy}\right)
 = \vcenter{\xy 0;/r1pc/:
\vtwistneg~{(-1,3.5)*+{\scriptstyle{a}}}{(0.5,3.5)*+{\scriptstyle{b}}}{(-1,2)}{(0.5,2)},
\vtwistneg~{(-1,2)}{(0.5,2)}{(-1,.5)}{(0.5,.5)},
(-.25,-.3)*{\bullet},
(-.25,-.3); (-1,.5) **\crv{(-1,.2)},
(-.25,-.3); (.5,.5) **\crv{(.5, .2)},
(-.25,-.3); (-.25,-1.2)*+{\scriptstyle{c}} **\dir{-},
(-.25,.1)*{\scriptstyle{\a}},
\endxy} =
\theta_a^{-1}\theta_b^{-1} \vcenter{\xy 0;/r1pc/:
\vcross~{(0,1)}{(.5,1)}{(0,0)}{(.5, 0)},
\vcross~{(.5,0)}{(1.5,0)}{(.5,-1)}{(1.5,-1.5)},
\vcross~{(0,-1)}{(.5,-1)}{(0,-2)}{(.5,-2)},
\vcross~{(-.5,0)}{(0,0)}{(-.5,-1)}{(0,-1)},
(-.5,-1);(-.5,-2) **\dir{-},
(-.5,0);(-.5,1) **\dir{-},
(-1.5,1)\vcap[1],
(-2,1)\vcap[2],
(.5,1);(.5,2)*+{\scriptstyle a} **\dir{-},
(1.5,0);(1.5,2)*+{\scriptstyle b} **\dir{-},
(-1.5,-2)\vcap[-1],
(.5,-2);(.5,-3) **\dir{-},
(-1.5,1);(-1.5,-2)**\dir{-},
(-2,-2)\vcap[-2],
(-2,1);(-2,-2) **\dir{-},
(1.5,-1.5);(1.5,-3) **\dir{-},
(1,-3.5)*{\bullet},
(1,-3.5); (.5,-3) **\crv{(.5,-3.2)},
(1,-3.5); (1.5,-3) **\crv{(1.5, -3.2)},
(1,-3.5);(1,-4.3)*+{\scriptstyle c} **\dir{-},
(1,-3)*{\scriptstyle{\a}},
\endxy} =
\theta_a^{-1}\theta_b^{-1} \, \vcenter{\xy 0;/r1pc/:
(0,2)\vcap[1],
\vtwistneg~{(1,2)}{(2,2)}{(1, 1)}{(2,1)},
(2,2); (2,3) **\dir{-},
(2,1); (2,-.5) **\dir{-}?(1)+(0.4,0)*{c},
(0,2); (0,1) **\dir{-},
(0,1)\vcap[-1],
(2,3)*{\bullet},(2,3.5)*{\scriptstyle{\a}},
(2,3);(1,4.5)*+{\scriptstyle a} **\dir{-},
(2,3);(3,4.5)*+{\scriptstyle b} **\dir{-},
\endxy}=
\frac{\theta_c}{\theta_a\theta_b} \vcenter{\xy 0;/r1pc/:
(0,0)*{\bullet},
(0,0); (-1,1.2)*+{\scriptstyle{a}} **\dir{-},
(0,0); (1,1.2)*+{\scriptstyle{b}} **\dir{-},
(0,0); (0,-1.2)*+{\scriptstyle{c}} **\dir{-},
(0,.5)*{\scriptstyle{\a}}
\endxy},
$$
where the second equality follows from \eqref{eq:ribbon2} and third equality is a consequence of the naturality of the ribbon structure. 
Thus, the matrix $[R_{a,b}^c]$ of $R_{a,b}^c$ relative to the chosen bases is $\frac{\theta_c}{\theta_a \theta_b} \id$.

Following from \eqref{eq:ribbon}, $(\theta_a \o \theta_a)\circ (R_{a,a})^2 = \theta_{a\o a}$ for all $a \in \Irr(\CC)$. So $(R_{a,a})^{2N}= \id_{a \o a}$ for all $a \in \Irr(\CC)$, where $N = \ord(\theta) = \mathop{\lcm}\limits_{c \in \Irr(\CC)} \ord(\theta_a)$. Thus, $R_{a,a}^c$ is a finite order operator on $\CC(a\o a,c)$ for any $c \in \Irr(\CC)$. In particular, $R_{a,a}^c$ is diagonalizable. Therefore, there exists a basis $B^{a,a}_c$ of $\CC(a \o a, c)$  consisting of eigenvectors of $R_{a,a}^c$. 

Let $\pi_\a = \vcenter{\xy 0;/r1pc/:
(0,0)*{\bullet},
(0,0); (-1,1.2)*+{\scriptstyle{a}} **\dir{-},
(0,0); (1,1.2)*+{\scriptstyle{a}} **\dir{-},
(0,0); (0,-1.2)*+{\scriptstyle{c}} **\dir{-},
(0,.5)*{\scriptstyle{\a}},
\endxy} \in B^{a,a}_c$ such that $R_{a,a}^c\left(
 \vcenter{\xy 0;/r1pc/:
(0,0)*{\bullet},
(0,0); (-1,1.2)*+{\scriptstyle{a}} **\dir{-},
(0,0); (1,1.2)*+{\scriptstyle{a}} **\dir{-},
(0,0); (0,-1.2)*+{\scriptstyle{c}} **\dir{-},
(0,.5)*{\scriptstyle{\a}},
\endxy}
\right) = \lambda_\a \,
 \vcenter{\xy 0;/r1pc/:
(0,0)*{\bullet},
(0,0); (-1,1.2)*+{\scriptstyle{a}} **\dir{-},
(0,0); (1,1.2)*+{\scriptstyle{a}} **\dir{-},
(0,0); (0,-1.2)*+{\scriptstyle{c}} **\dir{-},
(0,.5)*{\scriptstyle{\a}},
\endxy}$ 
for some $\lambda_\a \in \BC^\times$. Then, by the same calculation as above, we find
$$
\lambda_\a^2   \vcenter{\xy 0;/r1pc/:
(0,0)*{\bullet},
(0,0); (-1,1.2)*+{\scriptstyle{a}} **\dir{-},
(0,0); (1,1.2)*+{\scriptstyle{a}} **\dir{-},
(0,0); (0,-1.2)*+{\scriptstyle{c}} **\dir{-},
(0,.5)*{\scriptstyle{\a}},
\endxy} = (R_{a,a}^c)^2\left(
 \vcenter{\xy 0;/r1pc/:
(0,0)*{\bullet},
(0,0); (-1,1.2)*+{\scriptstyle{a}} **\dir{-},
(0,0); (1,1.2)*+{\scriptstyle{a}} **\dir{-},
(0,0); (0,-1.2)*+{\scriptstyle{c}} **\dir{-},
(0,.5)*{\scriptstyle{\a}},
\endxy}
\right)=\vcenter{\xy 0;/r1pc/:
\vtwistneg~{(-1,3.5)*+{\scriptstyle{a}}}{(0.5,3.5)*+{\scriptstyle{a}}}{(-1,2)}{(0.5,2)},
\vtwistneg~{(-1,2)}{(0.5,2)}{(-1,.5)}{(0.5,.5)},
(-.25,-.3)*{\bullet},
(-.25,-.3); (-1,.5) **\crv{(-1,.2)},
(-.25,-.3); (.5,.5) **\crv{(.5, .2)},
(-.25,-.3); (-.25,-1.2)*+{\scriptstyle{c}} **\dir{-},
(-.25,.1)*{\scriptstyle{\a}},
\endxy} =
\theta_a^{-2} \vcenter{\xy 0;/r1pc/:
\vcross~{(0,1)}{(.5,1)}{(0,0)}{(.5, 0)},
\vcross~{(.5,0)}{(1.5,0)}{(.5,-1)}{(1.5,-1.5)},
\vcross~{(0,-1)}{(.5,-1)}{(0,-2)}{(.5,-2)},
\vcross~{(-.5,0)}{(0,0)}{(-.5,-1)}{(0,-1)},
(-.5,-1);(-.5,-2) **\dir{-},
(-.5,0);(-.5,1) **\dir{-},
(-1.5,1)\vcap[1],
(-2,1)\vcap[2],
(.5,1);(.5,2)*+{\scriptstyle a} **\dir{-},
(1.5,0);(1.5,2)*+{\scriptstyle a} **\dir{-},
(-1.5,-2)\vcap[-1],
(.5,-2);(.5,-3) **\dir{-},
(-1.5,1);(-1.5,-2)**\dir{-},
(-2,-2)\vcap[-2],
(-2,1);(-2,-2) **\dir{-},
(1.5,-1.5);(1.5,-3) **\dir{-},
(1,-3.5)*{\bullet},
(1,-3.5); (.5,-3) **\crv{(.5,-3.2)},
(1,-3.5); (1.5,-3) **\crv{(1.5, -3.2)},
(1,-3.5);(1,-4.3)*+{\scriptstyle c} **\dir{-},
(1,-3)*{\scriptstyle{\a}},
\endxy}=
\frac{\theta_c}{\theta_a^2} \,\vcenter{\xy 0;/r1pc/:
(0,0)*{\bullet},
(0,0); (-1,1.2)*+{\scriptstyle{a}} **\dir{-},
(0,0); (1,1.2)*+{\scriptstyle{b}} **\dir{-},
(0,0); (0,-1.2)*+{\scriptstyle{c}} **\dir{-},
(0,.5)*{\scriptstyle{\a}},
\endxy}\,.
$$
Therefore, $\lambda_\a = \e_\a \frac{\sqrt{\theta_c}}{\theta_a}$ where $\e_\a = \pm 1$, and the matrix $[R_{a,a}^c]$ relative to the basis $B^{a,a}_c$ is of the form $\frac{\sqrt{\theta_c}}{\theta_a} \Delta_a^c$ where $\Delta_a^c$ is the matrix given by $(\Delta_a^c)_{\a, \a'} = \e_\a \delta_{\a, \a'}$.
\end{proof}
\begin{remark}\label{r1}
For any $a, c \in \Irr(\CC)$, the last proposition show that the eigenvalues of $R_{a,a}^c$ can only be $\pm \frac{\sqrt{\theta_c}}{\theta_a}$. Let $d_{\pm}$ be the multiplicity of the eigenvalue $\pm \frac{\sqrt{\theta_c}}{\theta_a}$ of $R_{a,a}^c$ respectively.  One can reorder the basis $B^{a,a}_c$ so that the matrix $[R_{a,a}^c]$ relative to the basis $B_c^{a,a}$ is given by
$[R_{a,a}^c]=   \frac{\sqrt{\theta_c}}{\theta_a}\Delta_a^c$ with
$$
\Delta_a^c = \left[\begin{array}{c|c} I_{d_+}& 0\\\hline 0& -I_{d_-} \end{array}\right]\,,
$$
where $I_d$ denotes the identity matrix of rank $d$
\end{remark}
The $S$ and $T$-matrices of a premodular category are insufficient to determine the matrix $[R_{a,a}^c]$ for the gauge choice in the preceding proposition as the multiplicity of 1 (or $-1$) in the diagonal of $\Delta_{a,c}$ cannot be determined.  However, we will prove that the $T$-matrix, the fusion coefficient $N^{a,a}_c$ and the \emph{generalized Frobenius-Schur indicator} $\nu^{\iota(c)}_{2,1}(a)$  determine $\Delta_a^c$, and hence $R_{a,a}^c$, up to a permutation of $B^{a,a}_c$, for all $a, c\in \Irr(\CC)$, where $\iota: \CC \to Z(\CC)$ is an embedding.  In particular, if $\CC$ is an MTC, $\nu^{\iota(c)}_{2,1}(a)$ and $N^{a,a}_c$ can be expressed in terms of the $S$- and $T$-matrices.
\section{The main result}
We continue to use the conventions developed in Section \ref{s1}, and consider a premodular subcategory  $\CC$ as a (premodular) subcategory of its center $Z(\CC)$. 

Recall that the objects in the center $Z(\CC)$ of $\CC$ are pairs $(x, \s_{x,-})$ in which $x \in \CC$ and $\s_{x,y}: x \o y \to y \o x$ is a half-braiding for any object $y$ of $\CC$ \cite{KasBook}. In particular, $\s_{x,y}$ is a natural isomorphism in $y$ for all $y \in \CC$. We will depict the half-braiding $\s_{x,y}$  of $(x, \s_{x,-})$  for any $y \in \CC$ as
$$
\s_{x,y} = \vcenter{\xy 0;/r.5pc/:
{\ar@[red]@{-}@`{(0,-2.5), (3,-3)} (0,0)*+{\scriptstyle \red{x}}; (3,-5)},
{\ar@{-}@`{(3,-2.5),(0,-3)}|(0.5){\hole} (3,0)*+{\scriptstyle y}; (0,-5)},
\endxy}\,.
$$
Note that the object $x$ is colored in red as the half-braiding $\s_{x, y}$ is only natural in $y$.

For $x \in \CC$, $\bx =(x, R^{-1}_{-,x}) \in Z(\CC)$. The assignment $x \mapsto \bx$ can be extended to a fully faithful braided tensor functor (or simply embedding) $\iota : \CC \to Z(\CC)$ which preserves the spherical structures  and hence the dimensions. So we can consider  $\CC$ as a premodular subcategory of $Z(\CC)$ under the identification $x \mapsto \bx = (x, R^{-1}_{-,x}) \in Z(\CC)$.  

Now we recall the definition \cite[\S 2]{NS4} of generalized Frobenius-Schur indicators $\nu_{2,1}^\bX(z)$ for any $\bX=(x, \s_{x,-}) \in Z(\DD)$ and $z \in \DD$, where $\DD$ is any \emph{strict spherical fusion category}. Consider the linear operator $E^{(2,1)}_{\bX, z}: \DD(x, z\o z) \to \DD(x, z\o z)$ given by
$$
E^{(2,1)}_{\bX, z}: \vcenter{\xy 0;/r1pc/:
(0,0)*{\bullet},
(0,0); (-1,-1.2)*+{\scriptstyle{z}} **\dir{-},
(0,0); (1,-1.2)*+{\scriptstyle{z}} **\dir{-},
{\ar@{-} (0,0);(0,1.2)*+{\scriptstyle x}},
\endxy} \mapsto \,\,
\vcenter{\xy 0;/r1pc/:
(0,0)*{\bullet},
{\ar@{-}@`{(-1, -1.2), (-2, 0),(0.2,1.7), (1,1.2)}|(0.55)\hole (0,0);(2.5,-1.2)*+{\scriptstyle z}},
{\ar@[red]@{-}@`{(0,1),(-1,.7)}(0,0);(-1,2)*+{\scriptstyle \red{x}}},
(0,0); (1,-1.2)*+{\scriptstyle{z}} **\dir{-},
\endxy}
$$
The superscript $(2,1)$ means there are 2 legs labeled by $z$, and the fisrt one is rotated to the other side of the second leg labelled by $z$ via the half-braiding $\s_{x,z}$.
The generalized Frobenius-Schur indicators $\nu_{2,1}^\bX(z)$ is defined as
$$
\nu_{2,1}^\bX(z) = \Tr(E^{(2,1)}_{\bX, z})\,.
$$
The readers are referred to  \cite[\S 2]{NS4} for more details on generalized Frobenius-Schur indicators, their properties and some applications.

Now we return to the embedding $\iota: \CC \to Z(\CC)$ described above for the premodular category $\CC$. In this case, for $c, a \in \Irr(\CC)$,  $\iota(c) = (c,\s_{c,-})$ with the half-braiding $\s_{c,a} = R_{a,c}^{-1}$. In terms of diagrams, we have
$$
\vcenter{\xy 0;/r.4pc/:
{\ar@[red]@{-}@`{(0,-2.5), (3,-3)} (0,0)*+{\scriptstyle \red{c}}; (3,-5)},
{\ar@{-}@`{(3,-2.5),(0,-3)}|(0.5){\hole} (3,0)*+{\scriptstyle a}; (0,-5)},
\endxy} = \vcenter{\xy 0;/r.4pc/:
{\ar@{-}@`{(0,-2.5), (3,-3)}|(0.5){\hole} (0,0)*+{\scriptstyle c}; (3,-5)},
{\ar@{-}@`{(3,-2.5),(0,-3)} (3,0)*+{\scriptstyle a}; (0,-5)},
\endxy}\,. 
$$  
Thus,
$$
E^{(2,1)}_{\iota(c), a}: \vcenter{\xy 0;/r1pc/:
(0,0)*{\bullet},
(0,0); (-1,-1.2)*+{\scriptstyle{a}} **\dir{-},
(0,0); (1,-1.2)*+{\scriptstyle{a}} **\dir{-},
{\ar@{-} (0,0);(0,1.2)*+{\scriptstyle c}},
\endxy} \mapsto \,\,
\vcenter{\xy 0;/r1pc/:
(0,0)*{\bullet},
{\ar@{-}@`{(-1, -1.2), (-2, 0),(0.2,1.7), (1,1.2)} (0,0);(2.5,-1.2)*+{\scriptstyle a}},
{\ar@{-}@`{(0,1),(-1,.7)}|(.55)\hole (0,0);(-1,2)*+{\scriptstyle c}},
(0,0); (1,-1.2)*+{\scriptstyle{a}} **\dir{-},
\endxy}
=\,
\vcenter{\xy 0;/r1pc/:
(0,0)*{\bullet},
{\ar@{-}@`{ (-1, -2),(-2, -1.5), (-1, .3)}|(.08)\hole (0,0);(1.5,-2.5)*+{\scriptstyle a}},
{\ar@{-}(0,0);(0,1.5)*+{\scriptstyle c}},
{\ar@{-}@`{ (2, -.4)}|(.72)\hole (0,0);(-.7,-2.5)*+{\scriptstyle a}},
\endxy}
=\,\theta_a\,
\vcenter{\xy 0;/r1pc/:
(0,0)*{\bullet},
{\ar@{-}(0,0);(0,1.5)*+{\scriptstyle c}},
{\ar@{-}@`{(.7,-.4)} (0,0);(.7,-1)},
{\ar@{-}@`{(-.7,-.4)} (0,0);(-.7,-1)},
\vtwistneg~{(-.7,-1)}{(.7,-1)}{(-.7,-2.5)}{(.7,-2.5)},
\endxy}\,.
$$
 Note that the semisimplicity of $\CC$ implies that the compositions of morphisms define a nondegenerate pairing 
 $$
 \CC(a \o b, c)\mathop{\o}\limits_\BC\CC(c, a \o b) \to \CC(c,c) =\BC
 $$
 for any $a,b, c \in \Irr(\CC)$. To compute the trace of $ E^{(2,1)}_{\iota(c), a}$, one can pick any gauge choice $B=\{B_c^{a,b} \mid a,b,c \in \Irr(\CC)\}$ of $\CC$. The pairing determines a dual basis $B^c_{a,b}$ for $\CC(c, a \o b)$ such that  
 $$
 \vcenter{\xy 0;/r1pc/:
(0,0)*{\bullet},
(0,-2)*{\bullet},
{\ar@{-}(0,0);(0,1.5)*+{\scriptstyle c}},
{\ar@{-}@`{(1,-.4), (1,-1.6)} (0,0);(0,-2)},
{\ar@{-}@`{(-1,-.4), (-1,-1.6)} (0,0);(0,-2)},
{\ar@{-}(0,-2);(0,-3.5)*+{\scriptstyle c}},
(-1.1, -1)*+{\scriptstyle a}, 
(1.1, -1)*+{\scriptstyle b}, 
(0, -1.6)*+{\scriptstyle \a}, 
(0, -.5)*+{\scriptstyle \b}, 
\endxy}=\,\delta_{\a, \b}
\vcenter{\xy 0;/r1pc/: 
{\ar@{-}(0,1.5)*+{\scriptstyle c};(0,-3.5)*+{\scriptstyle c}},
\endxy}\quad \text{ for}\quad \a, \b =1, \dots, N^{a,b}_c\,.
$$
Therefore, for any $a,c \in \Irr(\CC)$, 
\begin{equation} \label{eq:2nd_indicator_R}
\nu_{2,1}^{\iota(c)}(a) =\Tr(E^{(2,1)}_{\iota(c),a)})=\theta_a \sum_{\a =1}^{N^{a,a}_c} 
\vcenter{\xy 0;/r1pc/:
(0,0)*{\bullet},
{\ar@{-}(0,0);(0,1.5)*+{\scriptstyle c}},
{\ar@{-}@`{(.7,-.4)} (0,0);(.7,-1)},
{\ar@{-}@`{(-.7,-.4)} (0,0);(-.7,-1)},
\vtwistneg~{(-.7,-1)}{(.7,-1)}{(-.7,-2.5)}{(.7,-2.5)},
{\ar@{-}@`{(.7,-3.1)} (0,-3.5);(.7,-2.5)},
{\ar@{-}@`{(-.7,-3.1)} (0,-3.5);(-.7,-2.5)},
(0,-3.5)*{\bullet},
(0, -3.1)*+{\scriptstyle \a}, 
(0, -.4)*+{\scriptstyle \a}, 
{\ar@{-}(0,-3.5);(0,-5)*+{\scriptstyle c}},
\endxy} =\theta_a \Tr(R_{a,a}^c)\,. 
 \end{equation}
Hence $\Tr(R_{a,a}^c) =  \theta_a^{-1} \nu_{2,1}^{\iota(c)}(a)$.

It follows from Proposition \ref{p1} and the subsequent Remark \ref{r1} that $R_{a,a}^c$ can only have the eigenvalues $\frac{\pm \sqrt{\theta_c}}{\theta_a}$, and the matrix $[R_{a,a}^c]$ relative to some basis $B^{a,a}_c$ for $\CC(a \o a, c)$ is given by
$$
[R_{a,a}^c] =  \frac{\sqrt{\theta_c}}{\theta_a}\left[\begin{array}{c|c} I_{d_+}& 0\\\hline 0& -I_{d_-} \end{array}\right]
$$
where $d_\pm$ are the multiplicities of the eigenvalues $\frac{\pm \sqrt{\theta_c}}{\theta_a}$ respectively. Note that $d_++d_-= N^{a,a}_c$. By \eqref{eq:2nd_indicator_R},
$$
d_+- d_- = \frac{\theta_a}{\sqrt{\theta_c}} \Tr(R_{a,a}^c) =\frac{1}{\sqrt{\theta_c}} \nu_{2,1}^{\iota(c)}(a)
$$
and so
$$
0\le d_{\pm}= \frac{1}{2}(N^{a,a}_c \pm \frac{1}{\sqrt{\theta_c}} \nu_{2,1}^{\iota(c)}(a)).
$$
Thus,  $\frac{1}{\sqrt{\theta_c}} \nu_{2,1}^{\iota(c)}(a)$ is an integer which has the same parity as $N_c^{a,a}$ (i.e. they are congruent modulo 2).

The value of $\nu_{2,1}^{\iota(c)}(a)$ is completely determined by the modular data of the center $Z(\CC)$, and the forgetful functor $Z(\CC) \to \CC$ \cite[Cor.  5.6]{NS4}.  Precisely,  
\begin{equation}\label{eq:gen_2nd_indicator}
\nu_{2,1}^{\iota(c)}(a) = \frac{1}{\dim(\CC)} \sum_{\bX \in \Irr(Z(\CC))} [x: a]_\CC\,  \mathbb{S}_{\iota(c), \bX}  \, \bbtheta_\bX^2
\end{equation}
where $[x: a]_\CC = \dim \CC(x, a)$  if $\bX=(x, \sigma_{x,-})$, $\bbtheta_\bX$ and $\mathbb{S}_{\iota(c), \bX}$ are the modular data of $Z(\CC)$.    Since the value of $\Tr(R_{a,a}^c)$ and the fusion coefficient $N_c^{a,a}$ completely determine the multiplicities of the eigenvalues $\frac{\sqrt{\theta_c}}{\theta_a}$ of the matrix $[R_{a,a}^c]$ presented in Proposition \ref{p1} up to an ordering of the gauge choice. This proves the first part of following statement.
\begin{thm}\label{t:2}
Let $\CC$ be a premodular category over $\BC$. The braiding of $\CC$ is completely determined by its fusion rules, its ribbon structure and the generalized Frobenius-Schur indicators $\nu_{2,1}^{\iota(c)}(a)$ for all $a, c \in \Irr(\CC)$.  The value $\frac{1}{\sqrt{\theta_c}}\nu_{2,1}^{\iota(c)}(a)$ is an integer bounded by $N_c^{a,a}$ and has the same parity  as $N_c^{a,a}$ for any $a, c \in \Irr(\CC)$. 

The matrix $[R_{a,a}^c]$ in Proposition \ref{p1} is $\frac{\sqrt{\theta_c}}{\theta_a}\left[\begin{array}{c|c} I_{d_+}& 0\\\hline 0& I_{d_-} \end{array}\right]$ for some basis $B^{a,a}_c$ of $\CC(a\o a, c)$, and 
\begin{align}\label{eq:multipicity}
d_{\pm}= \frac{1}{2}(N^{a,a}_c \pm \frac{1}{\sqrt{\theta_c}} \nu_{2,1}^{\iota(c)}(a))\,.
\end{align}

In particular, if $\CC$ is modular, the braiding of $\CC$ is completely determined by the modular data $(S,T)$ of $\CC$ and
\begin{equation} \label{eq2}
    d_{\pm}
 =  \frac{1}{2}\left(N^{a,a}_c \pm \frac{1}{\sqrt{\theta_c} \dim \CC} \sum_{k, l \in \Irr(\CC)} d_k \ol{S}_{c,l} N^{kl}_a \frac{\theta^2_k}{\theta^2_l}\right)\,,
\end{equation}
where $d_k$ is the categorical dimension of the object $k$
\end{thm}
\begin{proof}
For the last assertion, it follows from \cite[Prop. 6.1]{NS4} that
\begin{equation}\label{eq:2nd_ind_modular}
\nu_{2,1}^{\iota(c)}(a) = \frac{1}{\dim \CC} \sum_{k, l \in \Irr(\CC)} d_k \ol{S}_{c,l} N^{kl}_a \frac{\theta^2_k}{\theta^2_l},
\end{equation}
where $S_{c,l}$ and $T_{c,l}=\delta_{c,l}\theta_l$ are the modular data of $\CC$. By the Verlinde formula, the fusion rules of $\CC$ are determined by the $S$-matrix. Therefore, the braiding of $\CC$ is completely determined by the modular data $(S,T)$ of $\CC$. 
\end{proof}
\begin{remark}  
Let $\CC$ be a modular tensor category. It was conjectured in \cite{PSS} that
$$
Y_{a,b}^c:=\frac{\sqrt{\theta_b}}{\sqrt{\theta_{c}}\dim \CC} \sum_{k, l \in \Irr(\CC)} \ol{S}_{b,k} \ol{S}_{c,l} N^{kl}_a \frac{\theta^2_k}{\theta^2_l} \in \BZ 
$$ 
for  all $a, b, c \in \Irr(\CC)$. It was further conjectured in \cite{BHS} that 
$$
\dim \CC(c, a\o a\o b) \pm Y_{a,b}^c \ge 0\,.
$$
Both conjectures were proved in \cite[\S 8]{NS4}, and it was further proved in \cite[Prop. 4.2]{BCT} that $\dim \CC(c, a\o a\o b) \pm Y_{a,b}^c$ is always an even integer. The second statement of Theorem \ref{t:2} could be considered as an extension of \cite[Prop. 4.2]{BCT} to premodular categories but only for $b = \1$.

The expression \eqref{eq:2nd_indicator_R} also appeared in \cite[Prop. 2.3 (1)]{BPD} with $\nu_{2,1}^{\iota(c)}(a)$ being substituted by \eqref{eq:2nd_ind_modular} and the Verlinde formula. Equation \eqref{eq2} can also be found in \cite{Bantay97}.
\end{remark}

\noindent
{\bf Acknowledgements:}
S.-H. N. was partially supported by NSF grant  DMS-1664418 and the Simons Foundations.
E.C.R. was partially supported by NSF grant DMS-2205962.  X.-G.W was partially
supported by NSF grant DMR-2022428 and by the Simons Collaboration on
Ultra-Quantum Matter, which is a grant from the Simons Foundation (651446,
XGW).  The authors have no relevant financial or non-financial interests to
disclose.

This material is based upon work supported by the National Science Foundation under Grant No. DMS-1928930, while the author was in residence at the Mathematical Sciences Research Institute in Berkeley, California, during the year of 2024.
\bibliographystyle{unsrt}
\bibliography{mybibl}

\end{document}